### v1

@@ 3.1. System Design
The technical architecture of Shiksha copilot (Figure 1) is grounded in the core design principles outlined earlier and comprises three components: Curriculum Ingestion, which converts textbooks into structured digital content for curriculum-aligned generation; Translation, a fine-tuned model for high-quality English–Kannada translation in the education domain; and the LLM Orchestrator, a modular agentic layer that coordinates tasks such as lesson planning, assessment generation, translation, and validation. These components are described in detail below.
@@ 3.2. System Implementation
In the Curation Phase, a team of 23 curators—15 English-medium and 8 Kannada-medium educators—used the web-based Curation Platform in Shiksha copilot to develop LPs. The process began with English-medium LPs: LLMs first generated draft Learning Objectives (LOs), which were reviewed and augmented by English-medium curators to ensure comprehensive curricular coverage and alignment with prescribed standards. These curated LOs then seeded LP generation via LLMs enhanced with ingested curriculum content through RAG. To account for the material constraints of target classrooms, prompts were designed to generate content suitable for low-resource settings.
@@
In the Distribution Phase, Shiksha copilot was deployed with 1,043 field teachers working in Kannada- and English-medium public schools across Karnataka, India. Most schools were low-resource, with limited digital infrastructure and a shortage of high-quality teaching and learning materials. Teachers accessed curated LPs via a web-based Distribution Platform, where they could create and customize their own copies.
@@
User Roles and Feature Access. Initially, all teachers were given standard user access. Standard users could view and manually customize curated LPs and use the AI-assisted Question Gen feature, which generated assessments aligned with curricular blueprints based on Bloom’s Taxonomy (Forehand, 2010). Based on platform engagement, 146 teachers were granted power user access. In addition to the standard features, power users could generate entirely new LPs using AI and apply AI-driven edits to curated LPs. They also had access to two interactive chatbot tools. Lesson Chat was automatically linked to each lesson plan, allowing teachers to ask lesson-specific questions such as concept clarifications or activity suggestions. Edu Chat, on the other hand, served as a general-purpose educational assistant for broader pedagogical support, including strategies for differentiated instruction.
@@ 4. Methods
Lesson Plan CurationTool Deployment with 1043 TeachersAug 2024Curation StartsOct 2024Curator Interviewsmid-Nov 2024Pre-survey DistributionDec 2024Tool Deployedmid-Jan 2025Teacher InterviewsMar 2025Post-survey Distribution
Figure 3. Study Timeline
@@ 4.2. Surveys
[…]
@@
[…]
@@ 4.3. Log Analysis
Curation Platform Logs. Log data from the Curation Platform was used to evaluate the quality of AI-generated LPs and to examine how curators engaged with and refined the content. A total of 991 curated LPs (965 English, 26 Kannada) were analyzed. For each lesson plan, we had access to both the original AI-generated version and the corresponding edited version produced by the curators. We utilized Python to computationally identify and quantify edits at the content block level across the LPs. Quantitative analysis focused on the frequency and distribution of edits across content blocks to uncover curation patterns. Curator ratings for each content block were also analyzed descriptively to assess perceptions of content quality. In parallel, a qualitative analysis was conducted using thematic analysis to categorize the types of edits made. The curators’ comments accompanying content block ratings were also coded thematically to extract insights into their evaluation rationale.
Distribution Platform Logs. Logs from the Distribution Platform were analyzed to understand how teachers interacted with curated LPs and AI-enabled components of the system. We examined a total of 5,544 LPs (4,064 English, 1,480 Kannada) created by the teachers using curated LPs. The edits by the teacher were computationally identified and analyzed using Python, following the same process as in the curation analysis. Quantitative analysis examined the number and frequency of edits per content block and per lesson plan. Thematic analysis was used to qualitatively examine the nature and instructional intent behind teacher modifications. Teacher ratings of LPs were also analyzed descriptively. Additionally, we analyzed teacher interactions with two chatbot features: Edu Chat and Lesson Chat. Quantitative metrics included the number of users, interaction frequency, and message volume. Chat transcripts were coded (see Appendix C) using thematic analysis to understand the nature of teacher–AI interactions and the types of support teachers sought from the system.
@@ 5.1. Collaborative Curation and Adaptation of AI-Generated Lesson Plans
The English-medium LPs were reported by curators to be of high quality and suitable for classroom use. Out of 7,744 content blocks reviewed, 96.13% were rated suitable for classroom use, 3.53% required minor adjustments, and only 0.34% were deemed unusable due to technical errors. These errors typically involved incomplete content generation or system-generated error messages and were addressed by regenerating the content. The quality of the LPs was closely tied to the quality of the Learning Objectives (LOs). Curators noted that the AI-generated LOs were often superficial and failed to capture the full breadth of the topic, requiring extensive edits. For example, in the Reflection of Light lesson, a Science curator observed that the LOs were limited to basic definitions, omitting key concepts such as image formation and mirror types. Curators revised the LOs to align with Bloom’s Taxonomy (Forehand, 2010), adding key terms to enhance conceptual depth and ensure comprehensive curricular alignment (see Appendix Table 3). They found that strengthening the LOs often led to high-quality LPs that required minimal subsequent editing. Curator 5, who curated 42 LPs, explained:
@@
Overall, edits to the English-medium LPs were minimal, with only 314 out of 7,744 content blocks (4.1%) requiring modifications. We categorized these edits into formatting edits (70.70%) and content edits (29.30%). Formatting edits involved minor adjustments such as line breaks, text styling (e.g., bold or italics), and punctuation corrections—none of which significantly altered the instructional content. Content edits addressed more substantive pedagogical issues, including clarifying explanations, ensuring curricular alignment, enhancing student engagement, and correcting factual inaccuracies. Assessment questions—particularly Multiple Choice Questions—were a common focus of content edits due to frequent inaccuracies in answer choices. Curators corrected these inaccuracies and annotated the correct answers clearly to aid teachers during classroom instruction. Curators exercised extra caution with assessment items due to their direct usage by students. Despite these focused corrections, the overall number of edits remained low.
@@
There was notable variability in edit distribution by subject area and curator. Subject-specific edit rates ranged from a low of 0.82% in Social Science to a high of 6.16% in English, indicating moderate variability (variance = 6.19, coefficient of variation = 0.75). English and Mathematics required more substantial edits compared to Social Science and Environmental Studies, where edits predominantly involved formatting. Variability among individual curators was substantially greater, with edit rates ranging widely from 0% to 16.67% ($\mu=4.87\%$, $\sigma=4.92$, variance = 24.22, CV = 1.01). A chi-square test confirmed significant differences among curators ($\chi^{2}$ = 272.39, df = 15, p ¡ 0.001), though with a small effect size (Cramer’s V = 0.1768). Notably, the variance attributed to individual curators was nearly four times greater than subject-specific variance (variance ratio = 3.91), suggesting individual reviewing styles significantly shaped editing practices more than differences across subjects. We also analyzed the ratings which curators gave to the generated LPs, and the Shannon entropy scores for the ratings, ranging from 0 (highly uniform ratings) to 1.35 (diverse use of rating categories), reinforced our finding that curators exhibited varying assessment patterns. Our analysis of interview data also supported these quantitative and log analyses, revealing that some curators perceived themselves as more rigorous reviewers compared to others. Nevertheless, despite variability at the curator level, the overall extent of editing remained consistently low, underscoring the general reliability of AI-generated content.
@@
While the edits in the English-medium content were minimal, curators made substantially more changes to the Kannada-medium LPs. Despite the pedagogical structure of the LPs being largely sound, curators identified widespread issues with the linguistic quality of the AI-translated Kannada content. Curator 14 who reviewed Kannada LPs explained:
@@
At the sentence level, curators observed that longer sentences often became grammatically incoherent. Literal translations from English frequently led to phrasing that was syntactically correct but semantically confusing or unnatural in Kannada. These sentence structures disrupted comprehension and required rewording to restore clarity. At the word level, the changes made were more nuanced and fell into several categories. First, curators replaced terms that were not typically used in Kannada-medium textbooks with standard equivalents. For example, svātantrya hōrāṭa (freedom fight) was replaced with svātantrya saṅgrāma (freedom struggle). Second, they corrected incorrect words that were slightly similar in meaning but inappropriate in context, such as guṇikānta being replaced with guṇākāra for multiplication, and samarūpate (uniformity) revised to samamiti (symmetry). Third, in cases where the meaning was accurate but not specific enough, curators substituted words that conveyed the intended meaning more precisely—for instance, kannaḍigaḷu (glasses) was replaced with darpaṇagaḷu (mirrors), and aṇṭugaḷu (glues) with aṇṭisuva sāmagrigaḷu (adhesive materials). Additional refinements were made to clarify vague or ambiguous terminology, such as replacing kālanukramaṇi (chronology) with kālarekhē (timeline), and vinimaya paddhati (exchange system) with vastuvinimaya paddhati (barter system).
@@
Our interviews found a more nuanced picture of engagement. While newer teachers tended to follow written plans more closely, more experienced teachers often relied on mental planning—drawing from professional experience and a range of informal resources. Many adapted or extended the plans using materials from trainings, peer networks, cluster meetings, and WhatsApp groups. However, since these adaptations were largely done mentally, they were rarely recorded in the system—especially due to time constraints. Teacher 7, who’s been teaching English for 18 years, shared:
@@ 5.2. How Teachers Used the AI System?
While the record-keeping practices varied—some school authorities accepted printed copies, while others still required handwritten ones—teachers consistently described Shiksha copilot as time-saving. It allowed them to either print the plans directly or use them as a reference when writing by hand. In both cases, it streamlined bureaucratic demands while freeing time for actual classroom preparation. Teacher 3 noted:
@@
Teachers utilized the AI-generated LPs extensively to support their classroom teaching. Although teachers already had access to other instructional resources through training programs, WhatsApp groups, online forums, and previous teaching materials, they valued Shiksha copilot’s LPs more because of the structured and detailed approach. The clearly articulated 5E framework significantly helped teachers organize instructional sequences effectively and provided a cohesive scaffold for lesson delivery. Teacher 15 with 20 years experience outlined the benefits of using AI-generated LPs:
@@
Teachers appreciated that the LPs included a wide range of classroom activity ideas, clear explanations of complex concepts, and contextually relevant examples. For instance, an early-career Grade 7 Math teacher highlighted how the Geometry lesson plan included relatable real-world examples, such as household objects and local buildings, making abstract concepts more accessible. Previously, teachers often relied on extensive external resources, including internet searches and multiple textbooks, to create engaging, activity-based lessons. Shiksha copilot significantly reduced this burden by providing comprehensive, ready-to-use content in one place. Teacher 18 with 10 years of teaching experience explained:
@@
Beyond LPs, teachers actively used Shiksha copilot’s chat features—Lesson Chat and Edu Chat—to quickly obtain supplemental teaching resources. Lesson Chat was used 111 times ($\mu=2.52$ messages/chat, $\sigma=2.43$), and Edu Chat was used 82 times ($\mu=1.95$ messages/chat, $\sigma=1.64$). Although Lesson Chat was intended for lesson-specific questions and Edu Chat for general educational queries, teachers often did not differentiate their use distinctly. Queries primarily focused on obtaining assessment questions for both formative and summative (Black, 1993) purposes, such as unit test papers, one-mark questions for science chapters, and objective-type questions on probability. Additionally, teachers requested engaging formative assessment activities, including puzzles or riddles for algebraic expressions, questions for agree and disagree activities for integers, and objective questions for slap the board activities. One teacher noted the efficiency gained by generating questions for the Slap the Board activity, which typically required over an hour but was completed in less than five minutes using Shiksha copilot. Teachers also sought practical classroom resources like phonics activities for AEIOU sounds, paper folding activities, and contextual information such as examples of crops in Karnataka, uses of forests in Kannada, and details about the Taj Mahal in Kannada.

### v2

@@ 3.1. System Design
The technical architecture of Guru is grounded in the core design principles outlined earlier and comprises three key components: Curriculum Ingestion, Translation, and LLM Orchestration.
@@ 3.2. System Implementation
In the Curation Phase, a team of 23 curators—15 English-medium and 8 Kannada-medium educators—used the web-based Curation Platform (as shown in Figure 3) in Shiksha Copilot to develop LPs. Curators were selected by the pedagogy team of the partner organization. They included a mix of experienced teachers and student educators with backgrounds in working with public school students. All curators held at least a bachelor’s degree in the relevant subject area for which they developed lesson plans. The process began with English-medium LPs: LLMs first generated draft Learning Objectives (LOs), which were reviewed and augmented by English-medium curators to ensure comprehensive curricular coverage and alignment with prescribed standards. These curated LOs then seeded LP generation via LLMs enhanced with ingested curriculum content through RAG. To account for the material constraints of target classrooms, prompts were designed to generate content suitable for low-resource settings.
Figure 3. Curator’s Platform for Editing LPs
@@
In the Distribution Phase, Shiksha Copilot was deployed with 1,043 field teachers working in Kannada- and English-medium public schools across Karnataka, India. All the schools were Kannada-medium, and a subset of them also included English-medium classrooms. Most schools were low-resource, with limited digital infrastructure and a shortage of high-quality teaching and learning materials. The partner organization collaborated with government officials to recruit teachers from 757 schools across 35 educational districts, ensuring statewide coverage. The officials played a critical role in identifying teachers for training and pilot implementation. Teachers were informed about the study prior to training and provided consent to participate through a pre-survey. They were also informed of their right to withdraw at any time, with clear instructions for doing so included in the consent form. Only the teachers who participated in the study had access to the tool. Teachers accessed curated LPs via a web-based Distribution (as shown in Figure 4, where they could create and customize their own copies.
Figure 4. LP View in the Teacher’s Platform
@@
User Roles and Feature Access. Initially, all teachers were given standard user access. Standard users could view and manually customize curated LPs and use the AI-assisted Question Gen feature, which generated assessments aligned with curricular blueprints based on Bloom’s Taxonomy (Forehand, 2010). Based on platform engagement, 146 teachers were granted power user access by the partner organization. This access was rolled out sequentially to gather structured feedback on the AI features from teachers and to build greater enthusiasm, drawing on the organization’s prior experience of working with them. In addition, the cost of AI usage was taken into account to ensure that the organization could plan for and manage expenses effectively based on actual usage patterns. In addition to the standard features, power users could generate entirely new LPs using AI and apply AI-driven edits to curated LPs. They also had access to two interactive chatbot tools. Lesson Chat was automatically linked to each lesson plan, allowing teachers to ask lesson-specific questions such as concept clarifications or activity suggestions. Edu Chat, on the other hand, served as a general-purpose educational assistant for broader pedagogical support, including strategies for differentiated instruction.
@@ 4. Methods
School TimelineStudy MilestonesLesson Plan CurationTool DeploymentJul2024Aug2024Sep2024Oct2024Nov2024Dec2024Jan2025Feb2025Mar2025Apr2025May2025Jun2025 First Term BeginsSecond Term BeginsExam Season BeginsSummer Break BeginsLP Curation BeginsCurator InterviewsPre-survey DistributionTeacher InterviewsPost-survey DistributionLog Analysis Data Cutoff
Figure 5. School & Study Timeline
@@ 4.2. Surveys
Teaching-Related Stress. Lesson planning is a major source of workload and stress for teachers (Mockler and Stacey, 2024; Rieg et al., 2007). To assess the tool’s impact on stress, we used the Teacher Stress Scale (TSS) developed by Chen et al. (Chen et al., 2022), which includes two subscales: stress arising from teaching demands and stress related to limited school-based support. The former assesses pressures such as excessive workload, limited preparation time, and challenges in addressing diverse student needs; the latter captures perceptions of insufficient administrative and peer support. To align with other instruments, we adapted the original 5-point Likert scale to a 7-point format (1 = “strongly disagree” to 7 = “strongly agree”). The subscales showed strong internal consistency in our data: teaching demands had $\alpha=.79$ (pre) and .83.83 (post), while school-based support had $\alpha=.70$ (pre) and .83.83 (post).
Access to Teaching Resources. We developed a custom scale to assess whether the tool enhanced teachers’ access to contextualized instructional resources. Items used a 7-point Likert format to evaluate ease of access to interactive activities, culturally relevant LPs, digital content, and real-world curriculum-aligned examples. Pedagogical experts from the implementation partner reviewed items for content validity. The scale demonstrated strong internal consistency ($\alpha=.80$ pre, .86.86 post).
@@
Trust in AI-Based Educational Technologies. We measured teachers’ trust in AI-based educational tools using an instrument adapted from Viberg et al. (Viberg et al., 2024), based on Mayer et al.’s trust framework (Mayer et al., 1995). The scale covers three domains: perceived benefits (e.g., efficiency, support), perceived concerns (e.g., reliability, transparency), and willingness to adopt AI. The scale had excellent internal consistency in our sample ($\alpha=.89$ pre, .91.91 post).
@@ 4.3. Log Analysis
Curation Platform Logs. Log data from the Curation Platform was used to evaluate the quality of AI-generated LPs and to examine how curators engaged with and refined the content. A total of 991 curated LPs (965 English, 26 Kannada) were analyzed. For each lesson plan, we had access to both the original AI-generated version and the corresponding edited version produced by the curators. We utilized Python to computationally identify and quantify edits at the content block level across the LPs. Quantitative analysis focused on the frequency and distribution of edits across content blocks to uncover curation patterns. Curator ratings for each content block were also analyzed descriptively to assess perceptions of content quality. In parallel, qualitative analysis was conducted using thematic analysis to categorize the types of edits made. We employed an inductive–deductive approach (Fuster-Guillén, 2023), iteratively reviewing curator edits and comments to identify emergent themes and then grouping them into higher-order categories. The curators’ comments accompanying content block ratings were coded at the content block level, with codes capturing both evaluative judgments and rationales for revisions.
Distribution Platform Logs. Logs from the Distribution Platform were analyzed to understand how teachers interacted with curated LPs and AI-enabled components of the system. We examined a total of 5,544 LPs (4,064 English, 1,480 Kannada) created by the teachers using curated LPs. Although the curated Kannada-medium LPs were fewer than the English-medium ones (as seen in the previous section), they were more extensively utilized for generating LPs for classroom use, as all teachers taught in Kannada. Kannada-medium teachers also relied on English-medium lesson plans when a corresponding Kannada version was unavailable, as they were able to understand English. The edits by the teacher were computationally identified and analyzed using Python, following the same process as in the curation analysis. Quantitative analysis examined the number and frequency of edits per content block and per lesson plan. Thematic analysis was used to qualitatively examine the nature and instructional intent behind teacher modifications. Teacher ratings of LPs were also analyzed descriptively. Additionally, we analyzed teacher interactions with two chatbot features: Edu Chat and Lesson Chat. Quantitative metrics included the number of users, interaction frequency, and message volume. Chat transcripts were thematically coded to capture recurrent patterns in teacher–AI interactions. We employed an inductive-deductive coding approach (Fuster-Guillén, 2023): first, transcripts were read iteratively to identify recurring themes, and then these were organized into categories (e.g., assessment generation, multilingual support, lesson plan requests) as reflected in the codebooks (Appendix C). Each transcript was segmented at the query level, and codes were assigned to individual queries rather than entire conversations.
@@ 5.1. Collaborative Curation and Adaptation of AI-Generated Lesson Plans
The English-medium LPs were reported by curators to be of high quality and suitable for classroom use. Out of 7,744 content blocks reviewed, 96.13% were rated suitable for classroom use, 3.53% required minor adjustments, and only 0.34% were deemed unusable due to technical errors. These errors typically involved incomplete content generation or system-generated error messages and were addressed by regenerating the content. The quality of the LPs was closely tied to the quality of the Learning Objectives (LOs). Curators noted that the AI-generated LOs were often superficial and failed to capture the full breadth of the topic, requiring extensive edits. For example, in the Reflection of Light lesson, a Science curator observed that the LOs were limited to basic definitions, omitting key concepts such as image formation and mirror types. Curators revised the LOs to align with Bloom’s Taxonomy (Forehand, 2010), adding key terms to enhance conceptual depth and ensure comprehensive curricular alignment (see Appendix Table 3). They found that strengthening the LOs often led to high-quality LPs that required minimal subsequent editing. Curator 5, who taught English and Social Science, and has curated 42 LPs, explained:
@@
Type of Edits. Overall, edits to the English-medium LPs were minimal, with only 314 out of 7,744 content blocks (4.1%) requiring modifications. We categorized these edits into formatting edits (70.70%) and content edits (29.30%). Formatting edits involved minor adjustments such as line breaks, text styling (e.g., bold or italics), and punctuation corrections—none of which significantly altered the instructional content. Content edits addressed more substantive pedagogical issues, including clarifying explanations, ensuring curricular alignment, enhancing student engagement, and correcting factual inaccuracies. Assessment questions—particularly Multiple Choice Questions—were a common focus of content edits due to frequent inaccuracies in answer choices. Curators corrected these inaccuracies and annotated the correct answers clearly to aid teachers during classroom instruction. Curators exercised extra caution with assessment items due to their direct usage by students. Despite these focused corrections, the overall number of edits remained low.
@@
Variation in Edits. There was notable variability in edit distribution by subject area and curator. Subject-specific edit rates ranged from a low of 0.82% in Social Science to a high of 6.16% in English, indicating moderate variability (variance = 6.19, coefficient of variation = 0.75). English and Mathematics required more substantial edits compared to Social Science and Environmental Studies, where edits predominantly involved formatting. The Mathematics curators reported that most of their edits addressed errors in the assessment section, such as incorrect options or repeated answers in multiple-choice questions—a pattern that was also confirmed through content analysis of the edits. In English, by contrast, the edits were largely subjective, suggesting that the edit rate depended more on the curator’s discretion. Overall, variability among individual curators was greater than differences across subjects, with edit rates ranging from 0% to 16.67% ($\mu=4.87\%$, $\sigma=4.92$, variance = 24.22, CV = 1.01). A chi-square test confirmed significant differences among curators ($\chi^{2}$ = 272.39, df = 15, p ¡ 0.001), though with a small effect size (Cramer’s V = 0.1768). Notably, the variance attributed to individual curators was nearly four times greater than subject-specific variance (variance ratio = 3.91), suggesting individual reviewing styles significantly shaped editing practices more than differences across subjects. We also analyzed the ratings which curators gave to the generated LPs, and the Shannon entropy scores for the ratings, ranging from 0 (highly uniform ratings) to 1.35 (diverse use of rating categories), reinforced our finding that curators exhibited varying assessment patterns. Our analysis of interview data also supported these quantitative and log analyses, revealing that some curators perceived themselves as more rigorous reviewers compared to others. Nevertheless, despite variability at the curator level, the overall extent of editing remained consistently low, underscoring the general reliability of AI-generated content.
@@
While the edits in the English-medium content were minimal, curators made substantially more changes to the Kannada-medium LPs. Despite the pedagogical structure of the LPs being largely sound, curators identified widespread issues with the linguistic quality of the AI-translated Kannada content. Curator 14 from a Science background who reviewed Kannada LPs explained:
@@
Nuanced Edits in Kannada. At the sentence level, curators observed that longer sentences often became grammatically incoherent. Literal translations from English frequently led to phrasing that was syntactically correct but semantically confusing or unnatural in Kannada. These sentence structures disrupted comprehension and required rewording to restore clarity. At the word level, the changes made were more nuanced and fell into several categories. First, curators replaced terms that were not typically used in Kannada-medium textbooks with standard equivalents. For example, svātantrya hōrāṭa (freedom fight) was replaced with svātantrya saṅgrāma (freedom struggle). Second, they corrected incorrect words that were slightly similar in meaning but inappropriate in context, such as guṇikānta being replaced with guṇākāra for multiplication, and samarūpate (uniformity) revised to samamiti (symmetry). Third, in cases where the meaning was accurate but not specific enough, curators substituted words that conveyed the intended meaning more precisely—for instance, kannaḍigaḷu (glasses) was replaced with darpaṇagaḷu (mirrors), and aṇṭugaḷu (glues) with aṇṭisuva sāmagrigaḷu (adhesive materials). Additional refinements were made to clarify vague or ambiguous terminology, such as replacing kālanukramaṇi (chronology) with kālarekhē (timeline), and vinimaya paddhati (exchange system) with vastuvinimaya paddhati (barter system).
@@
Our interviews found a more nuanced picture of engagement. While newer teachers tended to follow written plans more closely, more experienced teachers often relied on mental planning—drawing from professional experience and a range of informal resources. Teachers viewed Shiksha Copilot’s lesson plans as one of several resources they could access. Many adapted or extended the plans using materials from trainings, peer networks, cluster meetings, and WhatsApp groups. However, since these adaptations were largely done mentally, they were rarely recorded in the system—especially due to time constraints. Teacher 7, who’s been teaching English for 18 years, shared:
@@ 5.2. How Teachers Used the AI System?
While the record-keeping practices varied—some school authorities accepted printed copies, while others still required handwritten ones—teachers consistently described Shiksha Copilot as time-saving. It allowed them to either print the plans directly or use them as a reference when writing by hand. Authorities who insisted on handwritten lesson plans had long justified the practice on the grounds that writing by hand ensured teachers actively engaged with the content, rather than merely reproducing plans circulated online or within teacher groups. This requirement predated the advent of AI and reflected broader concerns about teacher accountability. Despite these differing requirements, teachers found that the tool streamlined bureaucratic demands and freed time for actual classroom preparation. Teacher 3, who has been teaching Mathematics for 17 years, noted:
@@
Teachers utilized the AI-generated LPs extensively to support their classroom teaching. Although teachers already had access to other instructional resources through training programs, WhatsApp groups, online forums, and previous teaching materials, they valued Shiksha Copilot’s LPs more because of the structured and detailed approach. The clearly articulated 5E framework significantly helped teachers organize instructional sequences effectively and provided a cohesive scaffold for lesson delivery. Teacher 15, a Mathematics teacher with 20 years of experience outlined the benefits of using AI-generated LPs:
@@
Teachers appreciated that the LPs included a wide range of classroom activity ideas, clear explanations of complex concepts, and contextually relevant examples. For instance, an early-career Grade 7 Math teacher highlighted how the Geometry lesson plan included relatable real-world examples, such as household objects and local buildings, making abstract concepts more accessible. Previously, teachers often relied on extensive external resources, including internet searches and multiple textbooks, to create engaging, activity-based lessons. Shiksha Copilot significantly reduced this burden by providing comprehensive, ready-to-use content in one place. Teacher 18, a Science teacher with 10 years of teaching experience, explained:
@@
Chat assistance for customized resources. Beyond LPs, teachers actively used Shiksha Copilot’s chat features—Lesson Chat and Edu Chat—to quickly obtain supplemental teaching resources. Lesson Chat was used 111 times ($\mu=2.52$ messages/chat, $\sigma=2.43$), and Edu Chat was used 82 times ($\mu=1.95$ messages/chat, $\sigma=1.64$). Although Lesson Chat was intended for lesson-specific questions and Edu Chat for general educational queries, teachers often did not differentiate their use distinctly. Queries primarily focused on obtaining assessment questions for both formative and summative (Black, 1993) purposes, such as unit test papers, one-mark questions for science chapters, and objective-type questions on probability. Additionally, teachers requested engaging formative assessment activities, including puzzles or riddles for algebraic expressions, questions for agree and disagree activities for integers, and objective questions for slap the board activities. One teacher noted the efficiency gained by generating questions for the Slap the Board activity, which typically required over an hour but was completed in less than five minutes using Shiksha Copilot. Teachers also sought practical classroom resources like phonics activities for AEIOU sounds, paper folding activities, and contextual information such as examples of crops in Karnataka, uses of forests in Kannada, and details about the Taj Mahal in Kannada.